OpenStudy (anonymous):

will medal please help -2(a+5)(a-6) simplify your answer

OpenStudy (anonymous):

2a -10 (a-6) 3a -16 i believe

OpenStudy (anonymous):

\[-2 \times a(a-6)+5(a-6)\]

OpenStudy (anonymous):

\[-2 \times( a ^{2}-6a+5a-30)\]

OpenStudy (anonymous):

\[-2 \times (a ^{2} -a -30)\]

OpenStudy (anonymous):

\[-2a ^{2} +2a +60\]

OpenStudy (anonymous):

help with a couple more

OpenStudy (anonymous):

just tag me :)